# hep-ph/***

### v1

@@ 4.2 Analytical considerations
with the limits for $x\gg 1$ being $F_{\gamma}(x)\approx-\frac{7}{12}-\frac{1}{6}\,\mbox{log}\,x$, $F_{Z}(x)\approx\frac{5}{2}-\frac{5}{2}\,\mbox{log}\,x$, $G_{Z}(x,0)=G_{Z}(0,x)\approx\frac{1}{2}\,\mbox{log}\,x$, $G_{Z}(x,x)\approx-\frac{x}{2}-\frac{x\,\mathrm{log}\,x}{1-x}$ and $F_{\mathrm{Xbox}}(x,x)\approx-\frac{x}{4}+\frac{1}{4}\,(13-6\,\mbox{log}\,x)$. Furthermore, the CR for $\mu-e$ conversion in nuclei is
@@ 4.3 Results for Case 1)
A dependence of the size of $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ on the angle $\theta_{R}$ is in general expected, since the matrix combination in square brackets in eq. (26) is not diagonal and the light neutrino mass spectrum is either given by eq. (38), (39) or (40). We first explore the impact of this free angle by fixing the two scales $\mu_{0}$ and $M_{0}$ to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. Light neutrino masses either follow NO or IO with the lightest neutrino mass being set to $m_{0}=0.03\,\mbox{eV}$ or $m_{0}=0.015\,\mbox{eV}$, respectively. The results for the BRs of $\mu\to e\gamma$ and $\mu\to 3\,e$ as well as the CR of $\mu-e$ conversion in aluminium are displayed in fig. 1 for different values of $\theta_{R}$, as indicated by the different colours, and for both light neutrino mass orderings in the left and right plots, respectively. The small vertical coloured bands in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ show the interval of $\theta_{L}$ which leads to an acceptable fit of the lepton mixing angles with $\chi^{2}\leq 27$, as defined in eq. (164) in appendix C. As mentioned, none of the current bounds on $\mu-e$ transitions can constrain the considered parameter space. Thus, we present in fig. 1 the exclusion due to the prospective limits as grey area, see eqs. (58) and (59); in darker and lighter grey in the case of Phase 1 and Phase 2 of the experiment Mu3E in the plots in the middle. The constraining power of the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is limited to the larger values of $\theta_{R}$ that are shown, while the bound on $\mathrm{BR}(\mu\to 3\,e)$ expected from Phase 1 of Mu3E can exclude all used values apart from vanishing $\theta_{R}$. Phase 2 of Mu3E and even more the prospective bound on $\mathrm{CR}(\mu-e,\mathrm{Al})$ have the potential to test all the displayed values of $\theta_{R}$ for $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. We note that for these scales the obtained BRs and CR are for most of the studied values of $\theta_{R}$ slightly larger for light neutrino masses with IO than for a NO light neutrino mass spectrum. We can estimate their size easily for $\theta_{R}=0$, since for this choice the matrix combination in square brackets in eq. (26) is automatically diagonal. We use that the BRs and CR are mainly proportional to $|\eta_{e\mu}|^{2}$, as argued in section 4.2. The quantity $\eta_{e\mu}$ is calculated from eq. (28). It turns out to depend on $\theta_{L}$ (which can be written in terms of $\sin\theta_{13}$, see [33]) as well as on the differences of the squares of the Yukawa couplings $\Delta y_{ij}^{2}$, $\Delta y_{ij}^{2}=y_{i}^{2}-y_{j}^{2}$,
@@
To comprehensively analyse the parameter space that we have scanned, we display in fig. 3 the results for the BRs and CR either as function of the scale $M_{0}$, $\mu_{0}$ or the free angle $\theta_{R}$. The different colour-coding indicates, on the one hand, which experimental constraint can/does exclude a data point, i.e. for grey points at least one of the prospective bounds on $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ is violated as well as the current limit on at least one of the elements of the matrix $\eta$, see eq. (54), while the orange points correspond to the situation in which at least one of the prospective bounds is violated, but all current constraints on the magnitude of the matrix elements $\eta_{\alpha\beta}$ are fulfilled and red points, eventually, represent data points that are excluded by the current limits on $\eta$, but not by the future bounds on the three studied $\mu-e$ transitions. [^10] We remark that not in all scans data points corresponding to red points are found. On the other hand, the different colours which can be read off from the colour bar signal the average value of the Yukawa coupling, $\overline{y}$, which is defined as
@@
Lastly, we display possible correlations among $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in fig. 4. Since all three observables mainly depend on $|\eta_{e\mu}|^{2}$, see section 4.2, we expect them to be proportional to each other. This is to a certain extent also visible in fig. 4. The shown results confirm the exclusion potential of the future experiments for the three different observables: limited for $\mathrm{BR}(\mu\to e\gamma)$ alone, considerable for $\mathrm{BR}(\mu\to 3\,e)$ and strongest for $\mathrm{CR}(\mu-e,\mathrm{Al})$. In particular, we see that more data points still compatible with the prospective limit on $\mathrm{BR}(\mu\to 3\,e)$ from Mu3E Phase 2 can be excluded by the bound expected from $\mu-e$ conversion in aluminium than vice versa, compare plot in the middle of fig. 4.
@@ 4.4 Results for Case 2)
Figure 5: Case 2), $t$ even ($u$ even). Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{u}{n}-\theta_{L}$-plane in the upper, middle and lower row. Left (right) plots are for light neutrino masses with NO (IO) and $m_{0}=0.03\,(0.015)\,\mathrm{eV}$. The scales $\mu_{0}$ and $M_{0}$ are set to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, respectively. Dark (light) blue and red regions indicate the regions of parameter space compatible at the $1\,\sigma$ ($3\,\sigma$) level with different prospective bounds on the BRs and CR, see text for details. In the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ grey regions indicate the parameter space in which the lepton mixing angles can be accommodated, i.e. dark (light) grey for $\chi^{2}\leq 100\,(300)$ and in the insets $\chi^{2}\leq 27$.
@@
As mentioned, the angle $\theta_{R}$ is irrelevant and thus not specified. For $\mathrm{BR}(\mu\to e\gamma)$, shown in the upper row, the dark (light) blue regions indicate the parameter space in which the prospective bound from the experiment MEG II is passed at the $1\,\sigma$ ($3\,\sigma$) level, see eq. (58). As one can see, the impact of this future limit is mild on the considered parameter space, while being slightly stronger for light neutrino masses with IO. Similarly, for $\mathrm{BR}(\mu\to 3\,e)$, see plots in the middle row, the dark (light) blue regions indicate the parameter space compatible with the expected bound from Mu3E Phase 1 at the $1\,\sigma$ ($3\,\sigma$) level, see eq. (58). We observe that about half of the parameter space in the shown plane can be disfavoured; again, for light neutrino masses with IO the exclusion reach is slightly larger. The impact of the prospective limit from Mu3E Phase 2 at the $1\,\sigma$ ($3\,\sigma$) level on the parameter space is displayed with the dark (light) red regions. We clearly see that the viable parameter space is considerably reduced. The lower row of fig. 5 contains the corresponding plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$. As expected, the future bounds from COMET and Mu2e, respectively, see eq. (59), can only be passed in small regions of the parameter space. These are shown in dark (light) blue for the bound from COMET at the $1\,\sigma$ ($3\,\sigma$) level and in dark (light) red for Mu2e at the same $\sigma$ levels. Since these bounds are very similar, the regions are largely super-imposed. As explained in section 3, the ratio $\frac{u}{n}$ and the free angle $\theta_{L}$ are also constrained by the requirement to accommodate the lepton mixing angles well. For this reason, we show in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ in dark (light) grey the parameter space which leads to $\chi^{2}\leq 100\,(300)$, when using the $\chi^{2}$-function as defined in eq. (164) in appendix C. These large values are chosen in order to ensure the visibility of the corresponding parameter space in the $\frac{u}{n}-\theta_{L}$-plane. In the insets the regions with $\chi^{2}\leq 27$ are displayed in grey. While compatible with the future bound from MEG II and mostly with the one from Mu3E Phase 1, this part of the parameter space will be excluded by Mu3E Phase 2 and the upcoming searches for $\mu-e$ conversion in aluminium for the chosen values of the two scales $\mu_{0}$ and $M_{0}$.
@@
In fig. 6 the corresponding results for $t$ odd ($u$ odd) are displayed for several fixed values of $\theta_{R}$. Again, $\mu_{0}$ and $M_{0}$ are set to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. As for $t$ even, we study the three different $\mu-e$ transitions for both light neutrino mass orderings, see left and right plots, respectively. In the upper row the parameter space which is compatible with the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ from MEG II at the $1\,\sigma$ level is shown in different colours corresponding to the six different values of $\theta_{R}$. As we can see, the smaller the shown value of $\theta_{R}$ is, the smaller the allowed parameter space in the $\frac{u}{n}-\theta_{L}$-plane is. Furthermore, the viable parameter space for light neutrino masses with IO is slightly reduced compared to the one for NO. The same observations can be made when analysing the parameter space allowed by the future limit on $\mathrm{BR}(\mu\to 3\,e)$ from Mu3E Phase 1 at the $1\,\sigma$ level. We note that, as expected, the resulting allowed portion of parameter space is considerably smaller than in the case of $\mathrm{BR}(\mu\to e\gamma)$. Applying the prospective bound on $\mu-e$ conversion in aluminium from COMET at the $1\,\sigma$ level to the parameter space, we find that only tiny regions remain allowed in the $\frac{u}{n}-\theta_{L}$-plane, see plots in the lower row in fig. 6. In these plots, we also display the areas in which the lepton mixing angles are accommodated to a certain degree, i.e. $\chi^{2}\leq 100$, see eq. (164) in appendix C, in different light colours (and with dashed contours) depending on the value of $\theta_{R}$. [^13] Again, a larger value of $\chi^{2}$ is used in order to increase visibility. In the insets, instead the parameter space with $\chi^{2}\leq 27$ is displayed. Overall, we see that it is very challenging, if not impossible, for all shown values of $\theta_{R}$ to reconcile the requirement to fit the lepton mixing angles well with passing the expected strong bounds on $\mu-e$ conversion in aluminium. We remark that for $\theta_{R}\approx\frac{\pi}{4}$ we obtain areas very similar those for $t$ even. This is consistent, since in this case the matrix combination in square brackets in eq. (26) is (nearly) diagonal and the light neutrino masses are (almost) proportional to the square of the Yukawa couplings, see eq. (30), as it always happens for $t$ even.
@@
For $t$ even, we take $s=1$ and $t=2$ (leading to $u=0$), as shown in fig. 7. This leads to $\sin^{2}\theta_{13}\approx 0.022$, $\sin^{2}\theta_{12}\approx 0.34$ and $\sin^{2}\theta_{23}=0.5$, in agreement with [33, 45, 47]. Clearly, there is no dependence on the angle $\theta_{R}$. Also the two different values of $\theta_{L}$ that permit a good fit to lepton mixing data lead to nearly identical results, since the solid and dashed curves are almost always super-imposed. Here the solid curve represents the results for the smaller value of the angle $\theta_{L}$, while the dashed curve corresponds to those obtained for the larger value of $\theta_{L}$; explicit values of $\theta_{L}$ can be found in e.g. [33, 47]. In fig. 7 we display the same experimental constraints for the BRs of $\mu\to e\,\gamma$ and $\mu\to 3\,e$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ as in fig. 2 for Case 1). The conclusions which can be drawn regarding the parameter space compatible with these future limits are very similar to those found for Case 1) and $\theta_{R}=0$, namely the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed for the shown interval of $M_{0}$, while the prospective limit from Mu3E Phase 1 can only constrain larger masses, i.e. $M_{0}\gtrsim 10\,(3)\,\mbox{TeV}$, for light neutrino masses with NO (IO). In contrast to this, Mu3E Phase 2 is expected to reduce the allowed parameter space to $M_{0}\lesssim 200\,\mbox{GeV}$ for light neutrino masses with NO and the future bound on $\mathrm{CR}(\mu-e,\mathrm{Al})$ forces $M_{0}$ to lie in the interval $5\,\mbox{TeV}\lesssim M_{0}\lesssim 7\,\mbox{TeV}$ for both light neutrino mass orderings – close to the value of $M_{0}$ which leads to a (complete) cancellation in the CR, see eq. (69).
@@
For $t$ odd, we employ as example $s=1$ and $t=1$ which give rise to $u=1$. For this value of $u$ and $n=14$, the lepton mixing angles read $\sin^{2}\theta_{13}\approx 0.022$, $\sin^{2}\theta_{12}\approx 0.34$ and $\sin^{2}\theta_{23}\approx 0.56$, see also [33, 45, 47]. Using the same conventions as in figs. 2 and 7, we display the results in fig. 8 for six different values of the angle $\theta_{R}$ in different colours. While for $t$ even, we hardly see any difference in the results for the two different values of $\theta_{L}$ that lead to a good fit of the lepton mixing angles, the solid and dashed curves can be clearly distinguished in fig. 8. We also observe that the larger of the shown values of $\theta_{R}$ result in smaller values of the BRs and CR, except for $\theta_{R}=\frac{\pi}{4}$. For $\theta_{R}=\frac{\pi}{4}$ we instead find very similar results as for the combination $s=1$, $t=2$ ($u=0$), see fig. 7. This is expected, since for $\theta_{R}=\frac{\pi}{4}$ the matrix combination in square brackets in eq. (26) is automatically diagonal and the light neutrino masses are proportional to the squares of the Yukawa couplings even for $t$ odd. Comparing the results for light neutrino masses with NO and IO, they appear to be very similar, compare plots in the left and the right in fig. 8. For the larger values of $\theta_{R}$ that we show the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed, while the expected limit from Mu3E Phase 1 is more constraining, especially for larger values of $M_{0}$. The future bound from Mu3E Phase 2 can exclude all parameter space for $\mu_{0}=1\,\mbox{keV}$, unless $M_{0}\lesssim 600\,(300)\,\mbox{GeV}$ for light neutrino masses with NO (IO) and certain larger values of $\theta_{R}$ are chosen. The limit expected from COMET (and also Mu2e), on the other hand, restricts the allowed parameter space to values of $M_{0}$ close to $M_{0}\approx 6.5\,\mbox{TeV}$, for which a cancellation occurs in the CR. Furthermore, we note that the choice $s=0$ and $t=1$, corresponding to $u=-1$, leads to similar results as shown in fig. 8.151515For this value of $u$ and $n=14$, the atmospheric mixing angle is best-fitted to $\sin^{2}\theta_{23}\approx 0.44$, while the reactor and the solar mixing angles are like for $u=0$ and $u=1$.

### v2

@@ 4.2 Analytical considerations
with the limits for $x\gg 1$ being $F_{\gamma}(x)\approx-\frac{7}{12}-\frac{1}{6}\,\mbox{log}\,x$, $F_{Z}(x)\approx\frac{5}{2}-\frac{5}{2}\,\mbox{log}\,x$, $G_{Z}(x,0)=G_{Z}(0,x)\approx\frac{1}{2}\,\mbox{log}\,x$, $G_{Z}(x,x)\approx-\frac{x}{2}-\frac{x\,\mathrm{log}\,x}{1-x}$ and $F_{\mathrm{Xbox}}(x,x)\approx-\frac{x}{4}+\frac{1}{4}\,(13-6\,\mbox{log}\,x)$. Furthermore, the CR for $\mu-e$ conversion in nuclei (N) is
@@ 4.3 Results for Case 1)
A dependence of the size of $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ on the angle $\theta_{R}$ is in general expected, since the matrix combination in square brackets in eq. (26) is not diagonal and the light neutrino mass spectrum is either given by eq. (38), (39) or (40). We first explore the impact of this free angle by fixing the two scales $\mu_{0}$ and $M_{0}$ to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. Light neutrino masses either follow NO or IO with the lightest neutrino mass being set to $m_{0}=0.03\,\mbox{eV}$ or $m_{0}=0.015\,\mbox{eV}$, respectively. The results for the BRs of $\mu\to e\gamma$ and $\mu\to 3\,e$ as well as the CR of $\mu-e$ conversion in aluminium are displayed in fig. 1 for different values of $\theta_{R}$, as indicated by the different colours, and for both light neutrino mass orderings in the left and right plots, respectively. The small vertical coloured bands in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ show the interval of $\theta_{L}$ which leads to an acceptable fit of the lepton mixing angles with $\chi^{2}\leq 27$, as defined in eq. (170) in appendix D. As mentioned, none of the current bounds on $\mu-e$ transitions can constrain the considered parameter space. Thus, we present in fig. 1 the exclusion due to the prospective limits as grey area, see eqs. (58) and (59); in darker and lighter grey in the case of Phase 1 and Phase 2 of the experiment Mu3E in the plots in the middle. The constraining power of the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is limited to the larger values of $\theta_{R}$ that are shown, while the bound on $\mathrm{BR}(\mu\to 3\,e)$ expected from Phase 1 of Mu3E can exclude all used values apart from vanishing $\theta_{R}$. Phase 2 of Mu3E and even more the prospective bound on $\mathrm{CR}(\mu-e,\mathrm{Al})$ have the potential to test all the displayed values of $\theta_{R}$ for $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. We note that for these scales the obtained BRs and CR are for most of the studied values of $\theta_{R}$ slightly larger for light neutrino masses with IO than for a NO light neutrino mass spectrum. We can estimate their size easily for $\theta_{R}=0$, since for this choice the matrix combination in square brackets in eq. (26) is automatically diagonal.
We use that the BRs and CR are mainly proportional to $|\eta_{e\mu}|^{2}$, as argued in section 4.2. The quantity $\eta_{e\mu}$ is calculated from eq. (28). It turns out to depend on $\theta_{L}$ (which can be written in terms of $\sin\theta_{13}$, see [36]) as well as on the differences of the squares of the Yukawa couplings $\Delta y_{ij}^{2}$, $\Delta y_{ij}^{2}=y_{i}^{2}-y_{j}^{2}$,
@@
To comprehensively analyse the parameter space that we have scanned, we display in fig. 3 the results for the BRs and CR either as function of the scale $M_{0}$, $\mu_{0}$ or the free angle $\theta_{R}$. [^12] Note that this plot can also be found in the proceedings [87]. The different colour-coding indicates, on the one hand, which experimental constraint can/does exclude a data point, i.e. for grey points at least one of the prospective bounds on $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ is violated as well as the current limit on at least one of the elements of the matrix $\eta$, see eq. (54), while the orange points correspond to the situation in which at least one of the prospective bounds is violated, but all current constraints on the magnitude of the matrix elements $\eta_{\alpha\beta}$ are fulfilled and red points, eventually, represent data points that are excluded by the current limits on $\eta$, but not by the future bounds on the three studied $\mu-e$ transitions. [^13] We remark that not in all scans data points corresponding to red points are found. On the other hand, the different colours which can be read off from the colour bar signal the average value of the Yukawa coupling, $\overline{y}$, which is defined as
@@
Lastly, we display possible correlations among $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in fig. 4. Since all three observables mainly depend on $|\eta_{e\mu}|^{2}$, see section 4.2, we expect them to be proportional to each other. This is to a certain extent also visible in fig. 4. The shown results confirm the exclusion potential of the future experiments for the three different observables: limited for $\mathrm{BR}(\mu\to e\gamma)$ alone, considerable for $\mathrm{BR}(\mu\to 3\,e)$ and even stronger for $\mathrm{CR}(\mu-e,\mathrm{Al})$.
@@ 4.4 Results for Case 2)
Figure 5: Case 2), $t$ even ($u$ even). Predictions for $\mathrm{BR}(\mu\to e\gamma)$, $\mathrm{BR}(\mu\to 3\,e)$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ in the $\frac{u}{n}-\theta_{L}$-plane in the upper, middle and lower row. Left (right) plots are for light neutrino masses with NO (IO) and $m_{0}=0.03\,(0.015)\,\mathrm{eV}$. The scales $\mu_{0}$ and $M_{0}$ are set to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$, respectively. Dark (light) blue and red regions indicate the regions of parameter space compatible at the $1\,\sigma$ ($3\,\sigma$) level with different prospective bounds on the BRs and CR, see text for details. In the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ grey regions indicate the parameter space in which the lepton mixing angles can be accommodated, i.e. dark (light) grey for $\chi^{2}\leq 100\,(300)$ and in the insets $\chi^{2}\leq 27$. Simplified versions of these plots, with only contour lines or only experimentally preferred regions, can be found in appendix E, see figs. 22 and 23.
@@
As mentioned, the angle $\theta_{R}$ is irrelevant and thus not specified. For $\mathrm{BR}(\mu\to e\gamma)$, shown in the upper row, the dark (light) blue regions indicate the parameter space in which the prospective bound from the experiment MEG II is passed at the $1\,\sigma$ ($3\,\sigma$) level, see eq. (58). As one can see, the impact of this future limit is mild on the considered parameter space, while being slightly stronger for light neutrino masses with IO. Similarly, for $\mathrm{BR}(\mu\to 3\,e)$, see plots in the middle row, the dark (light) blue regions indicate the parameter space compatible with the expected bound from Mu3E Phase 1 at the $1\,\sigma$ ($3\,\sigma$) level, see eq. (58). We observe that about half of the parameter space in the shown plane can be disfavoured; again, for light neutrino masses with IO the exclusion reach is slightly larger. The impact of the prospective limit from Mu3E Phase 2 at the $1\,\sigma$ ($3\,\sigma$) level on the parameter space is displayed with the dark (light) red regions. We clearly see that the viable parameter space is considerably reduced. The lower row of fig. 5 contains the corresponding plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$. As expected, the future bounds from COMET and Mu2e, respectively, see eq. (59), can only be passed in small regions of the parameter space. These are shown in dark (light) red for the bound from COMET at the $1\,\sigma$ ($3\,\sigma$) level and in dark (light) blue for Mu2e at the same $\sigma$ levels. Since these bounds are very similar, the regions are largely super-imposed. As explained in section 3, the ratio $\frac{u}{n}$ and the free angle $\theta_{L}$ are also constrained by the requirement to accommodate the lepton mixing angles well. For this reason, we show in the plots for $\mathrm{CR}(\mu-e,\mathrm{Al})$ in dark (light) grey the parameter space which leads to $\chi^{2}\leq 100\,(300)$, when using the $\chi^{2}$-function as defined in eq. (170) in appendix D. These large values are chosen in order to ensure the visibility of the corresponding parameter space in the $\frac{u}{n}-\theta_{L}$-plane. In the insets the regions with $\chi^{2}\leq 27$ are displayed in grey. While compatible with the future bound from MEG II and mostly with the one from Mu3E Phase 1, this part of the parameter space will be excluded by Mu3E Phase 2 and the upcoming searches for $\mu-e$ conversion in aluminium for the chosen values of the two scales $\mu_{0}$ and $M_{0}$.
@@
In fig. 6 the corresponding results for $t$ odd ($u$ odd) are displayed for several fixed values of $\theta_{R}$. Again, $\mu_{0}$ and $M_{0}$ are set to $\mu_{0}=1\,\mbox{keV}$ and $M_{0}=3\,\mbox{TeV}$. As for $t$ even, we study the three different $\mu-e$ transitions for both light neutrino mass orderings, see left and right plots, respectively. In the upper row the parameter space which is compatible with the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ from MEG II at the $1\,\sigma$ level is shown in different colours corresponding to the six different values of $\theta_{R}$. As we can see, the smaller the shown value of $\theta_{R}$ is, the smaller the allowed parameter space in the $\frac{u}{n}-\theta_{L}$-plane is. Furthermore, the viable parameter space for light neutrino masses with IO is slightly reduced compared to the one for NO. The same observations can be made when analysing the parameter space allowed by the future limit on $\mathrm{BR}(\mu\to 3\,e)$ from Mu3E Phase 1 at the $1\,\sigma$ level. We note that, as expected, the resulting allowed portion of parameter space is considerably smaller than in the case of $\mathrm{BR}(\mu\to e\gamma)$. Applying the prospective bound on $\mu-e$ conversion in aluminium from COMET at the $1\,\sigma$ level to the parameter space, we find that only tiny regions remain allowed in the $\frac{u}{n}-\theta_{L}$-plane, see plots in the lower row in fig. 6. In these plots, we also display the areas in which the lepton mixing angles are accommodated to a certain degree, i.e. $\chi^{2}\leq 100$, see eq. (170) in appendix D, in different light colours (and with dashed contours) depending on the value of $\theta_{R}$. [^16] Again, a larger value of $\chi^{2}$ is used in order to increase visibility. In the insets, instead the parameter space with $\chi^{2}\leq 27$ is displayed. Overall, we see that it is very challenging, if not impossible, for all shown values of $\theta_{R}$ to reconcile the requirement to fit the lepton mixing angles well with passing the expected strong bounds on $\mu-e$ conversion in aluminium. We remark that for $\theta_{R}\approx\frac{\pi}{4}$ we obtain areas very similar to those for $t$ even. This is consistent, since in this case the matrix combination in square brackets in eq. (26) is (nearly) diagonal and the light neutrino masses are (almost) proportional to the square of the Yukawa couplings, see eq. (30), as it always happens for $t$ even.
@@
For $t$ even, we take $s=1$ and $t=2$ (leading to $u=0$), as shown in fig. 7. This leads to $\sin^{2}\theta_{13}\approx 0.022$, $\sin^{2}\theta_{12}\approx 0.34$ and $\sin^{2}\theta_{23}=0.5$, in agreement with [36, 48, 52]. Clearly, there is no dependence on the angle $\theta_{R}$. Also the two different values of $\theta_{L}$ that permit a good fit to lepton mixing data lead to nearly identical results, since the solid and dashed curves are almost always super-imposed. Here the solid curve represents the results for the smaller value of the angle $\theta_{L}$ (which equals $\theta_{L}\approx 0.184$ for light neutrino masses of both orderings), while the dashed curve corresponds to those obtained for the larger value of $\theta_{L}$ (which sums up to $\pi$ together with the smaller value of $\theta_{L}$), compare [36, 52]. In fig. 7 we display the same experimental constraints for the BRs of $\mu\to e\,\gamma$ and $\mu\to 3\,e$ and $\mathrm{CR}(\mu-e,\mathrm{Al})$ as in fig. 2 for Case 1). The conclusions which can be drawn regarding the parameter space compatible with these future limits are very similar to those found for Case 1) and $\theta_{R}=0$, namely the future bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed for the shown interval of $M_{0}$, while the prospective limit from Mu3E Phase 1 can only constrain larger masses, i.e. $M_{0}\gtrsim 10\,(3)\,\mbox{TeV}$ for light neutrino masses with NO (IO). In contrast to this, Mu3E Phase 2 is expected to reduce the allowed parameter space to $M_{0}\lesssim 200\,\mbox{GeV}$ for light neutrino masses with NO and the future bound on $\mathrm{CR}(\mu-e,\mathrm{Al})$ forces $M_{0}$ to lie in the interval $5\,\mbox{TeV}\lesssim M_{0}\lesssim 8\,\mbox{TeV}$ for both light neutrino mass orderings – close to the value of $M_{0}$ which leads to a (complete) cancellation in the CR, see eq. (69).
@@
For $t$ odd, we employ as example $s=1$ and $t=1$ which give rise to $u=1$. For this value of $u$ and $n=14$, the lepton mixing angles read $\sin^{2}\theta_{13}\approx 0.022$, $\sin^{2}\theta_{12}\approx 0.34$ and $\sin^{2}\theta_{23}\approx 0.56$, see also [36, 48, 52]. Using the same conventions as in figs. 2 and 7, we display the results in fig. 8 for six different values of the angle $\theta_{R}$ in different colours. While for $t$ even, we hardly see any difference in the results for the two different values of $\theta_{L}$ that lead to a good fit of the lepton mixing angles, the solid and dashed curves can be clearly distinguished in fig. 8.181818The smaller value of $\overline{\theta}_{L}$, the angle appearing in the matrix $\tilde{U}_{\nu}$, see eq. (37), is given by $\overline{\theta}_{L}\approx 0.146$ for light neutrino masses with NO and $\overline{\theta}_{L}\approx 0.148$ for IO light neutrino masses, while the larger value always sums up to $\pi$ together with the smaller one. For completeness, we mention that for the choice $u=-1$ the same values of $\overline{\theta}_{L}$ are obtained. For further details, we refer to e.g. [36, 52]. We also observe that the larger of the shown values of $\theta_{R}$ result in smaller values of the BRs and CR, except for $\theta_{R}=\frac{\pi}{4}$. For $\theta_{R}=\frac{\pi}{4}$ we instead find very similar results as for the combination $s=1$, $t=2$ ($u=0$), see fig. 7. This is expected, since for $\theta_{R}=\frac{\pi}{4}$ the matrix combination in square brackets in eq. (26) is automatically diagonal and the light neutrino masses are proportional to the squares of the Yukawa couplings even for $t$ odd. Comparing the results for light neutrino masses with NO and IO, they appear to be very similar, compare plots in the left and the right in fig. 8. For the larger values of $\theta_{R}$ that we show the prospective bound on $\mathrm{BR}(\mu\to e\gamma)$ is passed, while the expected limit from Mu3E Phase 1 is more constraining, especially for larger values of $M_{0}$. The future bound from Mu3E Phase 2 can exclude all parameter space for $\mu_{0}=1\,\mbox{keV}$, unless $M_{0}\lesssim 600\,(300)\,\mbox{GeV}$ for light neutrino masses with NO (IO) and certain larger values of $\theta_{R}$ are chosen. The limit expected from COMET (and also Mu2e), on the other hand, restricts the allowed parameter space to values of $M_{0}$ close to $M_{0}\approx 6.5\,\mbox{TeV}$, for which a cancellation occurs in the CR. Furthermore, we note that the choice $s=0$ and $t=1$, corresponding to $u=-1$, leads to similar results as shown in fig. 8.191919For this value of $u$ and $n=14$, the atmospheric mixing angle is best-fitted to $\sin^{2}\theta_{23}\approx 0.44$, while the reactor and the solar mixing angles are like for $u=0$ and $u=1$.